OpenStudy (anonymous):

Hexachlorophene, a compound made up of atoms of carbon, hydrogen, chlorine, and oxygen, is an ingredient in germicidal soaps. Combustion of a 1.000 g sample yields 1.407 g of carbon dioxide, 0.134 g of water and 0.523 g of chlorine gas. What are the mass percents of carbon, hydrogen, oxygen and chlorine in hexachlorophene?

OpenStudy (anonymous):

Can someone explain to me how to answer it?

OpenStudy (abb0t):

mass of C in Hexachlorophene = mass of C in CO\(\sf _2\) mass of C in CO\(\sf _2\) = (1.407 g / 44 g/mol) x (1 mol C / 1 mol CO2) x 12 g/mol mass of C in CO2 = 0.3837 g = mass of C in Hexachlorophene % Cl = (0.3837 g / 1 g) x 100% = 38.37%=====> answer mass of H in Hexachlorophene = mass of H in H2O mass of H in H2O = (0.134 g / 18 g/mol) x (2 mol H / 1 mol H2O) x 1 g/mol mass of H in H2O = 0.0149 g = mass of H in Hexachlorophene % Cl = (0.0149 g / 1 g) x 100% = 1.49 %=====> answer mass of Cl in Hexachlorophene = mass of Cl in Cl2 gas mass of Cl in Cl2 gas = (0.523 g / 70.9 g/mol) x (2 mol Cl / 1 mol Cl2) x 35.45 g/mol mass of Cl in Cl2 gas = 0.523 g = mass of Cl in Hexachlorophene % Cl = (0.523 g / 1 g) x 100% = 52.3 %=====> answer % O in sample = 100% - 38.37% - 1.49% - 52.3% % O in sample = 7.84%

OpenStudy (anonymous):

Thank you so much!! Although how did you get the numbers to multiply it to get the mass?

OpenStudy (anonymous):

Nevermind. I found out they are conversion factors. Thank You!!